The Everton College paid $1862 for an order of 48 calculators. The store paid $11 for each scientific calculator The others all graphing calculators cost the store $57 each how many of each type of calculator was ordered? The store ordered ____ scientific and _____ graphing calculators.
Answer by checkley77(12844)

11s+57g=1,862
s+g=48 or s=48-g replace s with (48-g) in the equation above & solve for g.
11(48-g)+57g=1,862
528-11g+57g=1,862
46g=1,862-528
46g=1,334
g=1,334/46
g=29 graphing calculators were bought.
48-29=19 scientific calculators were bought.
Proof:
11*19+57*29=1,862
209+1,653=1,862
1,862=1,862